A study is done to see if there is a relationship between the cost of tuition of colleges and the average salary that graduates make 5 years after college.  A random sample of 100 colleges is taken and the tuition and average salary information is recorded for each college.  The key statistics are tabled below.

Cost of tuition

Mean = $17,500

SD = $3,700

Average salary of graduates

Mean = $33,000

SD = $9,000

Correlation between tuition and salary

r = .25

Sample size

n = 100

c.                   Predict the salary of a graduate from a college with a tuition of $25,000.
Answer #1

cost of tation - x = 17.500 , Sx = 3700 Average salary → V = 33000 , Sy-9000 8=0.25, n=100 slope (61) = (S/S) = (0.25) (9000/